what is the highest cycle efficiency possible for heat engine operating between 800 and 15
1
Expert's answer
2021-09-15T02:15:08-0400

Maximum efficiency is for engines.


For engines, we have


"T1\u200b=15+273=288K"


"T2\u200b=800+273=1073K"


"\u03b7=\\frac{T2\u200b\u2212T1}{T2}\u200b"


"\u03b7=\\frac{1073\u200b\u2212288}{1073}\u200b =0.7315"


"\u03b7 =0.7315"


Hence, 73.15%.